OpenStudy (anonymous):

confused ..do these two particles intersect? do they collide? r(t)sub1= and r(t)sub2=<1+2t,1+6t, 1+14t> @MIT 18.02 Multiva…

OpenStudy (zarkon):

their paths cross, but they don't occupy the same space at the same time

OpenStudy (anonymous):

what formula did you use

OpenStudy (zarkon):

I used the two vector valued functions you provided. no other formulas

OpenStudy (anonymous):

parametricequation route? p1=(0,0,0) v1=<1,1,1> p2=(1,1,1) v2=<2,6,4>

OpenStudy (zarkon):

in order to 'collide' they have to have the same 3 spatial coordinates at the same time....this is not possible

OpenStudy (zarkon):

both particles go through the points (1,1,1) and (2,4,8) but they do it at different times

OpenStudy (anonymous):

two meteors can't collide?

OpenStudy (anonymous):

ok, you want me to figure it out myself. Thank you

OpenStudy (zarkon):

Are they meteors now?

OpenStudy (across):

hmm

OpenStudy (across):

\[t=1+2t\]\[t^2=1+6t\]\[t^3=1+14t\]

OpenStudy (across):

the first equation implies\[t=-1\]

OpenStudy (across):

but then\[1\neq-5\]

OpenStudy (across):

nor\[-1\neq-13\]

OpenStudy (zarkon):

solve \[r_1(t)=r_2(s)\] you will find where they cross the same point. But s and t will be different... the way across wrote above is what I first did...then I checked to see if their paths ever crossed.

OpenStudy (anonymous):

thank yall